OpenStudy (anonymous):

What is 3x + 4y = 18 x + 3y = 11 Can you show me the process of how you got the answer. I'm only in 8th grade.

OpenStudy (anonymous):

(2, 3)

OpenStudy (anonymous):

\[x=11-3y => 3(11-3y) + 4y = 11 => 33-9y+4y = 11\] \[-9y+4y=11-33 => -5y= -22 => y = -22/-5 = 4.4\]

OpenStudy (anonymous):

Isolate y and do the same as I did to find its value.

OpenStudy (anonymous):

3x + 4y = 18 Multiply the second equa. for 3: 3x + 9y = 33 Subtract: 5y = 15 => y = 3 Replace: x = 11 - 3y = 11 - 3 * 3 = 2 Thus intersection point (2, 3)

OpenStudy (anonymous):

Wow, what a big mistake I made...

OpenStudy (anonymous):

Turned 3 into 33..

OpenStudy (anonymous):